# A Robust Mixed-Effects Bandit Algorithm for Assessing Mobile Health Interventions

### v3

@@ F.2 Preliminaries
By Lemma $11$ in Abbasi-Yadkori et al. (2011), we have
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\tilde{\sigma}_{i,t}^{2}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}\leq 2\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right).$
The lower bound on the weights implies
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}$ $\leq 2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right)$
as desired. Now,
$tr(\tilde{\sigma}_{i,t}^{2}\phi(x_{i,t})\phi(x_{i,t})^{\top})$ $=\tilde{\sigma}_{i,t}^{2}tr(\phi(x_{i,t})^{\top}\phi(x_{i,t}))$
$=\tilde{\sigma}_{i,t}^{2}\|\phi(x_{i,t})^{\top}\|^{2}$
$=3\sigma_{i,t}^{2}\|x_{i,t}\|^{2}$
$\leq\frac{3}{4}$
and $tr(L_{\otimes})=tr(L)\cdot tr(I)=KMd$. Then by Hadamard’s inequality we have
$\log\det(V_{K+1})$ $\leq 2Kd\log\left(\frac{1}{2Kd}\text{tr}\left(V_{K+1}\right)\right)$
$=2Kd\log\left(\frac{1}{2Kd}\left[\gamma 2Kd+\lambda 2KMd+\frac{3}{4}K\cdot(K+1)/2\right]\right)$
$=2Kd\log\left(\left[\gamma+\lambda M+3\frac{K+1}{16d}\right]\right)$
Finally, we state Azuma’s concentration inequality which describes concentration of super-martingales with bounded differences and is useful in controlling the regret due to the randomization of Thompson sampling.
Proposition 2 (Azuma’s concentration inequality).
If a super-martingale $(Y_{t})_{t\geq 0}$ corresponding to a filtration $\mathcal{F}_{t}$ satisfies $|Y_{t}-Y_{t-1}|<c_{t}$ some constant $c_{t}$ for all $t=1,\ldots,T$ then for any $\alpha>0$:
@@
$P\left(Y_{T}-Y_{0}\geq\alpha\right)\leq\exp\left(-\frac{\alpha^{2}}{2\sum_{t=1}^{T}c_{t}^{2}}\right).$
F.3 Proof of Theorem 1
We first decompose the regret bound
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $\pi_{i,t}^{\star}x(S_{i,t},a_{i,t}^{*})^{\top}\theta^{*}_{i,t}-\pi_{i,t}x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$
$=\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $(\pi_{i,t}^{\star}-\pi_{i,t})x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}$
$+\pi_{i,t}^{\star}\left\{x(S_{i,t},a_{i,t}^{*})^{\top}\theta^{*}_{i,t}-x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\right\}\big{]}$
$\leq\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $(\pi_{i,t}^{\star}-\pi_{i,t})x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$
$+\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}x(S_{i,t},a_{i,t}^{*})^{\top}\theta^{*}_{i,t}-x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$
@@ F.3 Proof of Theorem 1
For the first term, $\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}(\pi_{i,t}^{\star}-\pi_{i,t})x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$, one can apply a nearly identical proof to that of Greenewald et al. (2017). For the second term we need a novel strategy, and thus focus on that term. The proof follows closely from Abeille & Lazaric (2017) with several adjustments. Assumption 6 implies that we only need to consider the unit ball $\mathcal{X}=\{\|x\|\leq 1\}$. Then the second term of the regret can be decomposed into
@@
$\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left((\phi(x_{i,t}^{\star}))^{\top}\theta^{\star}-\phi(x_{i,t})^{\top}\tilde{\theta}_{k}\right)}_{R^{TS}(K)}+\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(\phi(x_{i,t})^{\top}\tilde{\theta}_{k}-\phi(x_{i,t})^{\top}\theta^{\star}\right)}_{R^{RLS}(K)}$
@@
where $\phi(x_{i,t}^{\star})$ is the context vector under the optimal action and $\theta^{\star}$ is the true parameter value. The first term is the regret due to the random deviations caused by sampling $\tilde{\theta}_{k}$ and whether it provides sufficient useful information about the true parameter $\theta^{\star}_{k}$. The second term is the concentration of the sampled term around the true linear model for the advantage function.
Definition 2.
We define the filtration $\mathcal{F}_{k}$ as the information accumulated up to stage $k$ before the sampling procedure, that is, $\mathcal{F}_{k}=(\mathcal{F}_{1},\sigma(x_{1},r_{2},x_{2},\ldots,x_{k-1},r_{k-1}))$, and filtration $\mathcal{F}_{k}^{x}$ as the information accumulated up to stage $k$ and including the sampled context, that is, $\mathcal{F}_{t}=(\mathcal{F}_{1},\sigma(x_{1},r_{2},x_{2},\ldots,x_{k-1},r_{k-1},x_{k}))$.
@@
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(\phi(x_{i,t})^{\top}\tilde{\theta}_{k}-\phi(x_{i,t})^{\top}\hat{\theta}_{k}\right)+\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(\phi(x_{i,t})^{\top}\hat{\theta}_{k}-\phi(x_{i,t})^{\top}\theta^{\star}_{k}\right)$
@@
The first term describes the deviation of the TS linear predictor from the RLS one, while the second term describes the deviation of the RLS linear predictor from the true linear predictor. The first term is controlled by the construction of the sampling distribution $D^{TS}$, while the second term is controlled by the RLS estimate being a minimizer of the regularized cumulative squared error in (LABEL:eq:dml_finalform). In particular, the first term will be small when the TS estimate concentrates around the RLS one, while the second will be small when the RLS estimate concentrates around the true parameter vector. The next proposition gives a lower bound on the probability that, for all stages, both the RLS parameter vector concentrates around the true parameter vector and the TS parameter vector concentrates around the RLS one.
Recall that
$\beta_{k}(\delta)=v_{k}\left[2\log\left(\frac{\det(V_{k})^{1/2}}{\det(V_{0})^{1/2}\delta/2}\right)\right]^{1/2}+B$
Proposition 3.
Let $\hat{E}_{k}$ denote the event that $\hat{\theta}_{k}$ concentrates around the true parameter for all $l\leq k$, i.e., $\hat{E}_{k}=\{\forall l\leq k,\|\hat{\theta}_{l}-\theta^{\star}_{l}\|_{V_{l}}\leq\beta_{l}(\delta^{\prime})\}$. Let $\gamma_{k}(\delta)\equiv\beta_{k}(\delta^{\prime})\sqrt{cd\log\frac{c^{\prime}d}{\delta}}$et $\tilde{E}_{k}$ denote the event that $\tilde{\theta}_{l}$ concentrates around the estimated parameter for all $l\leq k$, i.e., $\tilde{E}_{k}=\{\forall l\leq k,\|\tilde{\theta}_{l}-\hat{\theta}_{l}\|_{V_{l}}\leq\gamma_{l}(\delta^{\prime})\}$. Let $E_{k}=\hat{E}_{k}\cap\tilde{E}_{k}$. Then $P(E_{k})\geq 1-\delta/2$.
Proof.
Let $\delta^{\prime}=\delta/4K$, then Lemma 6 and a union bound give us
$P(\hat{E}_{K})$ $=P(\cap_{k=1}^{K}\{\|\hat{\theta}_{k}-\theta^{\star}\|_{V_{k}}\leq\beta_{k}(\delta^{\prime})\})$
$=1-\sum_{k=1}^{K}P(\|\hat{\theta}_{k}-\theta^{\star}\|_{V_{k}}>\beta_{k}(\delta^{\prime}))$
$=1-\sum_{k=1}^{K}\delta^{\prime}=1-\delta^{\prime}K=1-\delta/4.$
Applying the TS sampling distribution and $\tilde{\theta}_{k}=\hat{\theta}_{k}+\beta_{k}(\delta^{\prime})V_{k}^{-1/2}\eta_{k}$ where $\eta_{t}$ is drawn i.i.d. from $D^{TS}$ we have
$P\left(\|\tilde{\theta}_{k}-\hat{\theta}_{k}\|_{V_{k}}\leq\beta_{k}(\delta^{\prime})\sqrt{cd\log\left(\frac{c^{\prime}d}{\delta^{\prime}}\right)}\right)=P\left(\|\eta_{k}\|\leq\sqrt{cd\log\left(\frac{c^{\prime}d}{\delta^{\prime}}\right)}\right)\geq 1-\delta^{\prime}.$
by Definition 1. A union-bound argument yields the conclusion. ∎
We can then bound $R^{RLS}(K)$ by leveraging Lemma 6 and decomposing the error via
@@
$R^{RLS}(K)$ $\leq\sum_{k=1}^{K}\frac{1[E_{K}]}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|\phi(x_{i,t})^{\top}(\tilde{\theta}_{k}-\hat{\theta}_{k})|\right]$
$+\sum_{k=1}^{K}\frac{1[E_{K}]}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|\phi(x_{i,t})^{\top}(\hat{\theta}_{k}-\theta^{*})|\right]$
By definition of the event $E_{K}$, we have
@@
$|\phi(x_{i,t})^{\top}(\tilde{\theta}_{k}-\hat{\theta}_{k})|1[E_{k}]\leq\|\phi(x_{i,t})\|_{V_{k}^{-1}}\gamma_{k}(\delta^{\prime}),\quad|\phi(x_{i,t})^{\top}(\hat{\theta}_{k}-\theta^{\star})|1[E_{k}]\leq\|\phi(x_{i,t})\|_{V_{k}^{-1}}\beta_{k}(\delta^{\prime})$
so from Proposition 1, we have
$\sum_{k=1}^{K}\frac{1[E_{K}]}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|\phi(x_{i,t})^{\top}(\tilde{\theta}_{k}-\hat{\theta}_{k})|\right]$
$\leq\gamma_{K}(\delta^{\prime})\sum_{k=1}^{K}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\frac{1}{k}\|\phi(x_{i,t})\|_{V_{k}^{-1}}\right]$
@@
$\leq\gamma_{K}(\delta^{\prime})\sqrt{\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\frac{1}{k^{2}}}\sqrt{\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}}$
$\leq\gamma_{K}(\delta^{\prime})\sqrt{\sum_{k=1}^{K}\frac{1}{k}}\sqrt{\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}}$
$\leq\gamma_{K}(\delta^{\prime})\sqrt{H_{K}}\sqrt{\sum_{(i,t)\in\mathcal{O}_{k}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}}$
$\leq\gamma_{K}(\delta^{\prime})\sqrt{H_{K}}\sqrt{2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right)}.$
@@
Using a similar derivation for the $\beta_{k}(\delta^{\prime})$ case, we obtain
$R^{RLS}(K)$ $\leq\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\right)\sqrt{\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\frac{1}{k^{2}}}\sqrt{2c\log\left(\frac{\det\left(V_{K+1}\right)}{\det\left(V_{0}\right)}\right)}$
$\leq\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\right)\sqrt{\sum_{k=1}^{K}\frac{1}{k}}\sqrt{2c\log\left(\frac{\det\left(V_{K+1}\right)}{\det\left(V_{0}\right)}\right)}$
$\leq\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\right)\sqrt{H_{K}}\sqrt{2c\left[2Kd\log\left(\gamma+\lambda M+3\frac{K+1}{16d}\right)-\log\det\left(V_{0}\right)\right]}$
with probability at least $21-\delta/$ 2 by Proposition 3, where $H_{K}$ is the harmonic number. Note that $H_{K}\sim\log(K)$ for large $K$.

### v4

@@ E.4 Treatment Probability Estimation Errors
We begin by conditioning on $\mathcal{F}_{i,t}^{S}$ so that $x_{i,t}^{\top}\theta_{i,t}^{\star}$ is nonrandom. There are three cases.
Case 1: $x_{i,t}^{\top}\theta_{i,t}^{\star}=0$.
In this simple case, we have $(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}=0$.
Case 2: $x_{i,t}^{\top}\theta_{i,t}^{\star}>0$.
In this case, the optimal policy is to apply the control arm with probability $\pi_{i,t}^{\star}=\pi_{\min}$ because the treatment arm has a positive expected effect on the reward. Because $\pi_{i,t}\in[\pi_{\min},\pi_{\max}]$ by construction, we have $\pi_{i,t}\geq\pi_{\min}=\pi_{i,t}^{\star}$, which implies that $\pi_{i,t}-\pi_{i,t}^{\star}\geq 0$, proving the first statement in the proof for this case.
Now the probability $\pi_{i,t}$ is based on the following probability calculation:
$\Pr\{x_{i,t}^{{\top}}\tilde{\theta}_{i,t}<0\,|\,\mathcal{F}_{i,t}^{S}\}$
=\ = $\Pr\left[x_{i,t}^{{\top}}\left\{\theta_{i,t}^{\star}+(\check{\theta}_{i,t}-\theta_{i,t}^{\star})+\beta_{i,t}(\delta_{1},\delta_{2})\underline{\hbox{\set@color$V$}}_{i,t}^{-1/2}\eta_{i,t}\right\}<0\,\Big{|}\,\mathcal{F}_{i,t}^{S}\right]$
=\ = $\Pr\left[x_{i,t}^{{\top}}\left\{(\check{\theta}_{i,t}-\theta_{i,t}^{\star})+\beta_{i,t}(\delta_{1},\delta_{2})\underline{\hbox{\set@color$V$}}_{i,t}^{-1/2}\eta_{i,t}\right\}<-x_{i,t}^{{\top}}\theta_{i,t}^{\star}\,\Big{|}\,\mathcal{F}_{i,t}^{S}\right]$
[…]
=\ = $\Pr\left[\left\|(\check{\theta}_{i,t}-\theta_{i,t}^{\star})+\beta_{i,t}(\delta_{1},\delta_{2})\underline{\hbox{\set@color$V$}}_{i,t}^{-1/2}\eta_{i,t}\right\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\geq\frac{x_{i,t}^{{\top}}\theta_{i,t}^{\star}}{\left\|x_{i,t}\right\|_{\underline{\hbox{\set@color$V$}}_{i,t}^{-1}}}\,\Bigg{|}\,\mathcal{F}_{i,t}^{S}\right]$
[…]
where the last two lines follow from Markov’s inequality and Assumption 9. Working directly with the expectation, we can show that
$\mathbb{E}\left(\left\|(\check{\theta}_{i,t}-\theta_{i,t}^{\star})+\beta_{i,t}(\delta_{1},\delta_{2})\underline{\hbox{\set@color$V$}}_{i,t}^{-1/2}\eta_{i,t}\right\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}\,\Big{|}\,\mathcal{F}_{i,t}^{S}\right)$
[…]
=\ = $\left\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\right\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}+\beta^{2}_{i,t}(\delta_{1},\delta_{2})\mathbb{E}\left\|\eta_{i,t}\right\|^{2}.$
This result implies that
$\Pr\{x_{i,t}^{{\top}}\tilde{\theta}_{i,t}<0\}\leq\frac{\left\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\right\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}+\beta^{2}_{i,t}(\delta_{1},\delta_{2})\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{(x_{i,t}^{{\top}}\theta_{i,t}^{\star})^{2}\alpha\min(i,t)}\eqcolon M_{i,t}.$
Provided $M_{i,t}\leq\pi_{\min}$, we have $(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}=0$ because $\pi_{i,t}=\pi_{i,t}^{\star}=\pi_{\min}$. Alternatively, when $M_{i,t}>\pi_{\min}$, we have
$x_{i,t}^{{\top}}\theta_{i,t}^{\star}<\sqrt{\frac{\left\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\right\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}+\beta^{2}_{i,t}(\delta_{1},\delta_{2})\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{\pi_{\min}\alpha\min(i,t)}}.$
Thus, in both cases we have
$(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}<\sqrt{\frac{\left\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\right\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}+\beta^{2}_{i,t}(\delta_{1},\delta_{2})\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{\pi_{\min}\alpha\min(i,t)}}.$
Case 3:
This case is essentially identical to Case 2, except both $\pi_{i,t}-\pi_{i,t}^{\star}$ and $x_{i,t}^{\top}\theta_{i,t}^{\star}$ are negative. The resulting bound is
$(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}<\sqrt{\frac{\left\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\right\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}+\beta^{2}_{i,t}(\delta_{1},\delta_{2})\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{(1-\pi_{\max})\alpha\min(i,t)}}.$
Conclusion:
Combining the three cases yields
$(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}<\sqrt{\frac{\left\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\right\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}+\beta^{2}_{i,t}(\delta_{1},\delta_{2})\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{\min(\pi_{\min},1-\pi_{\max})\alpha\min(i,t)}}.$
Lemma 13.
Under event $E_{K}$, the following bound holds for all $(i,t)\in\mathcal{O}_{i,t}$:
$(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}<\frac{\tau\beta_{K}(\delta_{1},\delta_{2})}{\sqrt{\min(i,t)}},\quad\tau\coloneq\sqrt{\frac{1+\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{\alpha\min(\pi_{\min},1-\pi_{\max})}}.$
Proof.
Let $(i^{\star},t^{\star})$ denote the final decision point at stage $K$ and let $\delta_{1}^{\prime}=\delta_{1}/\{K(K+1)\}$. Conditional on $\mathcal{F}^{S}_{i^{\star},t^{\star}}$, Lemma 12 implies that
$(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}<\sqrt{\frac{\left\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\right\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}+\beta^{2}_{i,t}(\delta_{1}^{\prime},\delta_{2})\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{\min(\pi_{\min},1-\pi_{\max})\alpha\min(i,t)}}$
for all $(i,t)\in\mathcal{O}_{K}$ because $\mathcal{F}^{S}_{i,t}\subseteq\mathcal{F}^{S}_{i^{\star},t^{\star}}$. We can then apply Lemma 11 to argue that
@@ Conclusion:
$(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}<$ $\sqrt{\frac{\beta^{2}_{K}(\delta_{1},\delta_{2})+\beta^{2}_{K}(\delta_{1},\delta_{2})\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{\min(\pi_{\min},1-\pi_{\max})\alpha\min(i,t)}}$
=\ = $\frac{\beta_{K}(\delta_{1},\delta_{2})}{\sqrt{\min(i,t)}}\sqrt{\frac{1+\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{\alpha\min(\pi_{\min},1-\pi_{\max})}}$
under event $E_{K}$ for all $(i,t)\in\mathcal{O}_{i,t}$. ∎
E.5 Proof of Theorem 1
We first decompose the regret
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $\pi_{i,t}^{\star}(\bar{A}_{i,t}^{\star}|S_{i,t})x(S_{i,t},\bar{A}_{i,t}^{\star})^{\top}\theta_{i,t}^{\star}-\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})x(S_{i,t},\bar{A}_{i,t})^{\top}\theta_{i,t}^{\star}\big{]}$
$=\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$$\{\pi_{i,t}^{\star}(\bar{A}_{i,t}^{\star}|S_{i,t})-\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})\}x(S_{i,t},\bar{A}_{i,t})^{\top}\theta_{i,t}^{\star}$
$+\pi_{i,t}^{\star}(\bar{A}_{i,t}^{\star}|S_{i,t})\left\{x(S_{i,t},\bar{A}_{i,t}^{\star})^{\top}\theta_{i,t}^{\star}-x(S_{i,t},\bar{A}_{i,t})^{\top}\theta_{i,t}^{\star}\right\}\big{]}$
$\leq\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$$\{\pi_{i,t}^{\star}(\bar{A}_{i,t}^{\star}|S_{i,t})-\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})\}x(S_{i,t},\bar{A}_{i,t})^{\top}\theta_{i,t}^{\star}\big{]}$
$+\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{\{}x(S_{i,t},\bar{A}_{i,t}^{\star})^{\top}\theta_{i,t}^{\star}-x(S_{i,t},\bar{A}_{i,t})^{\top}\theta_{i,t}^{\star}\big{\}}.$
We can bound the (absolute value of) the first term using Lemma 13:
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}\{\pi_{i,t}^{\star}(\bar{A}_{i,t}^{\star}|S_{i,t})-\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})\}x(S_{i,t},\bar{A}_{i,t})^{\top}\theta_{i,t}^{\star}\big{]}$
<\ < $\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\frac{\tau\beta_{K}(\delta_{1},\delta_{2})}{\sqrt{\min(i,t)}}$
[…]
=\ = $\tau\beta_{K}(\delta_{1},\delta_{2})\sum_{k=1}^{K}\frac{2\sqrt{2}}{\sqrt{k}}$
[…]
=\ = $\tau\beta_{K}(\delta_{1},\delta_{2})4\sqrt{2K},$
under event $E_{K}$.
@@ E.5 Proof of Theorem 1
The proof for the second term follows closely from Abeille & Lazaric (2017) with several adjustments. Assumption 5 implies that we only need to consider the unit ball $\mathcal{X}=\{x\in\mathbb{R}^{d}:\|x\|\leq 1\}$. Then the second term of the regret can be decomposed into
@@
$\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(x_{i,t}^{\star\top}\theta_{i,t}^{\star}-x_{i,t}^{\top}\tilde{\theta}_{i,t}\right)}_{R^{TS}(K)}+\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(x_{i,t}^{\top}\tilde{\theta}_{i,t}-x_{i,t}^{\top}\theta_{i,t}^{\star}\right)}_{R^{RLS}(K)}$
@@
where $x_{i,t}^{\star}=x(S_{i,t},\bar{A}_{i,t}^{\star})$ and $x_{i,t}=x(S_{i,t},\bar{A}_{i,t})$. The first term is the regret due to the random deviations caused by sampling $\tilde{\theta}_{i,t}$ and whether it provides sufficient useful information about the true parameter $\theta_{i,t}^{\star}$. The second term is the concentration of the sampled term around the true linear model for the advantage function.
@@
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(x_{i,t}^{\top}\tilde{\theta}_{i,t}-x_{i,t}^{\top}\check{\theta}_{i,t}\right)+\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(x_{i,t}^{\top}\check{\theta}_{i,t}-x_{i,t}^{\top}\theta_{i,t}^{\star}\right)$
@@
The first term describes the deviation of the TS linear predictor from the RLS one, while the second term describes the deviation of the RLS linear predictor from the true linear predictor. The first term is controlled by the construction of the sampling distribution $\mathcal{D}^{TS}$, while the second term is controlled by the RLS estimate being a minimizer of the regularized cumulative squared error in (3). In particular, the first term will be small when the TS estimate concentrates around the RLS one, while the second will be small when the RLS estimate concentrates around the true parameter vector. Under event $E_{K}$, we can bound $R^{RLS}(K)$ by leveraging Lemma 6 and decomposing the error via
@@
$R^{RLS}(K)\leq\sum_{k=1}^{K}\frac{1}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|x_{i,t}^{\top}(\tilde{\theta}_{i,t}-\check{\theta}_{i,t})|\right]+\sum_{k=1}^{K}\frac{1}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|x_{i,t}^{\top}(\check{\theta}_{i,t}-\theta_{i,t}^{\star})|\right].$
Under event  $E_{K}$, we have
@@
$|x_{i,t}^{\top}(\tilde{\theta}_{i,t}-\check{\theta}_{i,t})|\leq\|x_{i,t}\|_{\underline{\hbox{\set@color$V$}}_{i,t}^{-1}}\gamma_{K}(\delta_{1},\delta_{2}),\quad|x_{i,t}^{\top}(\check{\theta}_{i,t}-\theta_{i,t}^{\star})|\leq\|x_{i,t}\|_{\underline{\hbox{\set@color$V$}}_{i,t}^{-1}}\beta_{K}(\delta_{1},\delta_{2}),$
where $\gamma_{K}(\delta_{1},\delta_{2})\coloneq\beta_{K}(\delta_{1},\delta_{2})\sqrt{cd\log\{K(K+1)c^{\prime}d/\delta_{1}\}}$ and $\beta_{K}(\delta_{1},\delta_{2})$ is defined in Lemma 11. Then, from Assumption 9, we have
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|x_{i,t}^{\top}(\tilde{\theta}_{i,t}-\check{\theta}_{i,t})|$
[…]
@@
=\ = $\gamma_{K}(\delta_{1},\delta_{2})\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|x_{i,t}\|_{\underline{\hbox{\set@color$V$}}_{i,t}^{-1}}$
<\ < $\gamma_{K}(\delta_{1},\delta_{2})\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\frac{1}{\sqrt{\alpha\min(i,t)}}$
[…]
=\ = $\frac{\gamma_{K}(\delta_{1},\delta_{2})}{\sqrt{\alpha}}\sum_{k=1}^{K}2\sqrt{\frac{2}{k}}$
[…]
=\ = $\gamma_{K}(\delta_{1},\delta_{2})4\sqrt{2K/\alpha}$
@@
under event $E_{K}$. Using a similar derivation for the $\beta_{K}(\delta_{1},\delta_{2})$ case, the following inequality holds under event $E_{K}$:
$R^{RLS}(K)$ <$<\ \left\{\beta_{K}(\delta_{1},\delta_{2})+\gamma_{K}(\delta_{1},\delta_{2})\right\}4\sqrt{2K/\alpha}$
$=\ \beta_{K}(\delta_{1},\delta_{2})\left[1+\sqrt{cd\log\{K(K+1)c^{\prime}d/\delta_{1}\}}\right]4\sqrt{2K/\alpha}$